The product of two consecutive integers is 9 less than 9 times their sum. Find the integers. I would like to know how to do this problem
Answer by stanbon(75887)

The product of two consecutive integers is 9 less than 9 times their sum. Find the integers.
----------------
1st: x
2nd: x+1
------
Equation:
x(x+1) = 9(x + x+1)+9
-------
x^2 + x = 9(2x+1)+9
x^2 + x = 18x+18
------------
x^2 -17x -18 = 0
(x-18)(x+1) = 0
Positive solution:
x = 18
x+1 = 19
=============
Cheers,
Stan H.